OpenStudy (anonymous):

What is the inverse of f(x)=2x+12

OpenStudy (anonymous):

to find the inverse just solve for x. f(x) = 2x+12 ---> f(x)-12= 2x ---> x = (f(x)-12)/2 For formalities you should switch the 'f(x)' and 'x' at this point for the correct answer of f(x) = (x-12)/2.

OpenStudy (anonymous):

y=2x+12 x = 2y+12 y = (x-12)/2 \[y=\frac{1}{2}x-6\]